# Boltzmanngasse 5, A 1090 Vienna, Austria

### v2

@@ 3.8.2 Closed or almost closed causal curves, time machine
$-\bar{W}(\bar{V}-2\bar{m}\bar{r})\prod_{i=1}^{N}\frac{(\bar{\mu}^{i})^{2}}{\bar{\Xi}_{i}}=\det\hat{g}=\pm N^{2}\det h=\frac{\det h}{\hat{g}(\hat{\nabla}\bar{t},\hat{\nabla}\bar{t})}=\frac{\det h}{\bar{g}(\bar{\nabla}\bar{t},\bar{\nabla}\bar{t})}\,.$ (3.60)
So ¯$\bar{\nabla}\bar{t}$ is timelike at $p\in\{\bar{V}-2\bar{m}\bar{r}>0\}$ if and only if the matrix $h$ defined in (3.58) is positive definite at $p$. Alternatively, one can compare (3.59) with (3.50) to reach the same conclusion.
Let ${\mycal N}\approx\mathbb{T}^{N}$ denote any principal orbit of the action of the isometry group generated by the Killing vectors fields $\partial_{\phi^{i}}$, and consider the pullback, which we denote by $h_{\mycal N}$, of $h=h_{ij}d\phi^{i}d\phi^{j}$ to ${\mycal N}$. When non-degenerate, $h_{\mycal N}$ is the metric induced by $\bar{g}$ on ${\mycal N}$. Since $h_{\mycal N}$ is Riemannian on the set $\{\bar{V}-2\bar{m}\bar{r}<0\}$, which is stably causal as ¯$\bar{\nabla}\bar{r}$ is timelike there, we have proved:
@@
1. All connected components of the set $\{\bar{V}-2\bar{m}\bar{r}<0\}$ are stably causal, with ¯$\bar{\nabla}\bar{r}$ timelike there.
@@
and then ¯$\bar{\nabla}\bar{t}$ is timelike there.
@@ 3.9.1 Negative Ricci curvature
$\ -\frac{\ell^{-2}(1-\ell^{-2}\bar{r}^{2})}{\bar{W}}\bigg{[}|\bar{\mu}|^{4}\big{(}\sum_{i=1}^{N}\frac{\bar{a}_{i}n^{i}dn^{i}}{\bar{\Xi}_{i}}\big{)}^{2}+2|\bar{\mu}|^{3}(\sum_{i=1}^{N}\frac{\bar{a}_{i}(n^{i})^{2}d|\bar{\mu}|}{\bar{\Xi}_{i}})(\sum_{j=1}^{N}\frac{\bar{a}_{j}n^{j}dn^{j}}{\bar{\Xi}_{i}})\bigg{]}$
@@ 3.10 Black hole regions
We claim that $({\mycal M}_{\mathrm{KS}},\bar{g})$ contains a black hole region or a white hole region, depending upon the choice of time orientation. To see this, recall that the gradient of $\bar{r}$ is a time function on the set $\{\bar{V}-2\bar{m}\bar{r}<0\}\cap\{\bar{r}<\bar{r}_{H}\}$. Let us choose the time orientation so that $-r$ is a time function on $\{\bar{V}-2\bar{m}\bar{r}<0\}$ near $\bar{r}_{H}$, thus $\bar{r}$ is strictly decreasing along every causal curve in this last region. Hence no causal curve that intersects
@@ 4 Projection diagrams
$\breve{g}:=$ $\bar{W}(1-\ell^{-2}\bar{r}^{2})d\bar{t}^{2}+\frac{\bar{U}d\bar{r}^{2}}{\bar{V}-2\bar{m}\bar{r}}+\frac{2\bar{m}\bar{r}}{\bar{U}}\biggl{(}\bar{W}d\bar{t}+\sum_{i=1}^{N}\frac{\bar{a}_{i}(\bar{\mu}^{i})^{2}d\phi^{i}}{\bar{\Xi}_{i}}\biggr{)}^{2}$
@@
with ±$\pm N^{2}$ given by (3.50), thus $\breve{\gamma}$ is obtained by ignoring the $h$-part of $\hat{g}$. It follows from (4.2) and (3.47) that, in the region where $h$ is positive definite we have, for any vector $X$,
@@
In order to extract out of $\mathring{\gamma}\equiv\breve{U}^{-1}\breve{\gamma}$ a two-dimensional Lorentzian metric which depends only upon $r$, we define
@@ 4.1 $\bar{V}-2\bar{m}\bar{r}<0$
±$\pm N^{2}=$ $-\frac{\bar{W}(\bar{V}-2\bar{m}\bar{r})}{\left(1+\frac{2\bar{m}\bar{r}}{\bar{U}}\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{(\bar{r}^{2}+\bar{a}_{i}^{2})\bar{\Xi}_{i}}\right)\prod_{i=1}^{N}(\bar{r}^{2}+\bar{a}_{i}^{2})}\,,$ (4.9)
@@
$\mathring{\gamma}(X,X)=$ $\breve{U}^{-1}N^{2}(X^{\bar{t}})^{2}+\frac{\bar{F}^{2}}{\bar{V}-2\bar{m}\bar{r}}(X^{\bar{r}})^{2}$
@@ 4.2 $\bar{V}-2\bar{m}\bar{r}>0$
On this region the coordinate $\bar{r}$ has spacelike gradient, which implies that the tensor field $\gamma$ of (3.47) has Lorentzian signature. On the subset where the matrix $h$ of (3.58) is positive definite we have ±$\pm N^{2}<0$, hence we need to choose the negative sign in the left-hand side of (4.9). Then
$\mathring{\gamma}(X,X)=$ $-\breve{U}^{-1}N^{2}(X^{\bar{t}})^{2}+\frac{\bar{F}^{2}}{\bar{V}-2\bar{m}\bar{r}}(X^{\bar{r}})^{2}$
@@ 4.3 Further extensions
Figure 4.1: The projection diagram for an extended spacetime with $\bar{m}>0$, in which there are no closed causal curves, with $\bar{V}-2\bar{m}\bar{r}$ having two zeros $-\ell<\bar{r}_{-}<\ell<\bar{r}_{+}$. Both horizons are of bifurcate-type, with non-zero surface gravity. The light blue regions are where $\nabla\bar{r}$ is timelike and the pink ones where $\nabla\bar{t}$ is timelike. The solid-blue curves represent the level sets of constant $\bar{r}$ while the dashed-red those of constant $\bar{t}$. The Kerr–Schild extension is indicated by the regions marked I-III. The diagram can be continued indefinitely in the vertical direction so that a shift of the figure by two diamonds generates a $\mathbbm{Z}$-group of isometries. Distinct maximal analytic extensions can be obtained by quotienting the spacetime by distinct subgroups of $\mathbbm{Z}$, leading to closed timelike curves through every point. Further distinct maximal analytic extensions can be obtained by removing, e.g., one or more bifurcation surfaces and passing to covering spaces.
Figure 4.2: The projection diagram for an extended spacetime with $\bar{m}>0$ in which there are no closed causal curves and $\bar{V}-2\bar{m}\bar{r}$ has three zeros $-\ell<\bar{r}_{1}<\bar{r}_{2}<\ell<\bar{r}_{3}$, with the zero at $\bar{r}_{2}$ of order two. The coloring of diamonds and the dotting of curves are as in Figure 4.1. The extension can be continued in all directions. The horizons at $\bar{r}_{2}$ are degenerate, in the sense that they have vanishing surface gravity $\kappa$, the other ones have non-zero $\kappa$.
Figure 4.3: The projection diagram for an extended spacetime with $\bar{m}>0$ in which there are no closed causal curves and $\bar{V}-2\bar{m}\bar{r}$ has four zeros $-\ell<\bar{r}_{1}<\bar{r}_{2}<\bar{r}_{3}<\ell<\bar{r}_{4}$. The extension can be continued to the whole Minkowskian plane in all directions. All horizons are of bifurcate-type, with non-zero surface gravity. The nature of the diamonds and of the curves are as in Figure 4.1. The Kerr–Schild extension is the union of the regions marked I to V. The extension is symmetric with respect to reflections across vertical and horizontal lines passing through the bifurcation points of the Killing horizons. Shifts of the plane by two diamonds along the vertical axis and by two diamonds along the horizontal axis generate a $\mathbbm{Z}^{2}$ group of isometries of the depicted spacetime when extended to the whole of $\mathbb{R}^{2}$; taking quotients by distinct subgroups of $\mathbbm{Z}^{2}$ leads to distinct, non-isometric, analytic extensions of the original metric, possibly but not necessarily introducing causality violations. Further distinct extensions can be obtained by passing to distinct coverings of the projection plane.
@@
Figure 4.4: Projection diagrams for the Kerr-Newman – anti-de Sitter metrics with two distinct zeros of the Boyer-Lindquist function $\Delta_{r}$ (left diagram) and one double zero (right diagram), from COS , generated using SzybkaOelz . The arrows represent the orientation of the orbits of the Killing vector $\partial_{t}$. The shaded area is the projection of the Carter time-machine region, including the singular ring, and should be removed from the diagram for a faithful representation of causality. An identical diagram is obtained for our metrics in the two-zeros case of Figure 4.1 when causality violations occur with $\bar{m}>0$.
@@ 5 Odd dimensions? Positive $\Lambda$?
$t$ $=i\bar{t}\,,\quad r=-i\bar{r}\,,\quad\,m=i^{-(d-1)}\bar{m}\,,\quad a_{i}=-i\bar{a}_{i}\,,\quad 1\leq i<N\,,\qquad$ (5.1)
@@ Appendix A Determinant
$h\equiv h_{ij}d\phi^{i}d\phi^{j}=$ $\underbrace{\sum_{i=1}^{N}\frac{\bar{r}^{2}+\bar{a}_{i}^{2}}{\bar{\Xi}_{i}}(\bar{\mu}^{i})^{2}(d\phi^{i})^{2}}_{A}+\frac{2\bar{m}\bar{r}}{\bar{U}}\underbrace{\biggl{(}\sum_{i=1}^{N}\frac{\bar{a}_{i}(\bar{\mu}^{i})^{2}d\phi^{i}}{\bar{\Xi}_{i}}\biggr{)}^{2}}_{uu^{T}}\,,$ (A.7)
$N^{j}=$ $h^{ij}\frac{2\bar{m}\bar{r}\bar{W}\bar{a}_{i}(\bar{\mu}^{i})^{2}}{\bar{U}\bar{\Xi}_{i}}\,,$ (A.8)
±$\pm N^{2}=$ $\bar{W}(1-\ell^{-2}\bar{r}^{2})+\frac{2\bar{m}\bar{r}}{\bar{U}}\bar{W}^{2}-h_{ij}N^{i}N^{j}\,.$ (A.9)
Next, there is the standard decomposition of the determinant of $\hat{g}$ in terms of ±$\pm N^{2}$ and $h$;
@@
Now to calculate ±$\pm N^{2}$ we need to know $N^{i}$ and hence we need also $h^{ij}$. Assuming from now on that we are away from zeros of $\left(1+\frac{2\bar{m}\bar{r}}{\bar{U}}\,u^{T}A^{-1}u\right)$, using (A.2) we obtain
@@
Now we can obtain ±$\pm N^{2}$ from (A.9);
±$\pm N^{2}$ $=\bar{W}(1-\ell^{-2}\bar{r}^{2})+\frac{2\bar{m}\bar{r}}{\bar{U}}\bar{W}^{2}\left(1-\frac{2\bar{m}\bar{r}}{\bar{U}}\frac{u^{T}A^{-1}u}{1+\frac{2\bar{m}\bar{r}}{\bar{U}}\,u^{T}A^{-1}u}\right)$
@@
±$\pm N^{2}$ $=\frac{\bar{W}}{1+\frac{2\bar{m}\bar{r}}{\bar{U}}\,u^{T}A^{-1}u}\left((1-\ell^{-2}\bar{r}^{2})+\frac{2\bar{m}\bar{r}}{\bar{U}}\left[1+\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{\bar{r}^{2}+\bar{a}_{i}^{2}}\right]\right)$
@@
Note that we have recovered (A) without dividing by $1-\ell^{-2}\bar{r}^{2}$. However, in this alternative calculation we have instead a division-with-cancellation by $\left(1+\frac{2\bar{m}\bar{r}}{\bar{U}}\,u^{T}A^{-1}u\right)$, which vanishes at the boundary of the causality-violating regions. Since the previous calculation did not necessitate such a division, and since the zero-sets do not overlap in the fully-rotating case by (3.39), we conclude that (A.18) holds without the restrictions mentioned. Hence::
@@
Steps 2-4 are taken care since they only set up the problem. In particular, since we did not change the coordinates $\bar{\mu}^{i}$ for $1\leq i\leq N$ we have simply $\tilde{\gamma}_{ij}=\bar{g}_{ij}$. Thus, the only thing left to calculate is
@@
To calculate the inverse metric we use (A.2). Plugging (A.2) into (A.28) gives (note that the expression $1+u\cdot D^{-1}\cdot u$ appearing below has already been calculated in (A))

### v3

@@ 3.8.2 Closed or almost closed causal curves, time machine
$\bar{W}(\bar{V}-2\bar{m}\bar{r})\prod_{i=1}^{N}\frac{(\bar{\mu}^{i})^{2}}{\bar{\Xi}_{i}}=\det\hat{g}=\pm N^{2}\det h=\frac{\det h}{\hat{g}(\hat{\nabla}\bar{t},\hat{\nabla}\bar{t})}=\frac{\det h}{\bar{g}(\bar{\nabla}\bar{t},\bar{\nabla}\bar{t})}\,.$ (3.60)
So $\bar{\nabla}\bar{t}$ is timelike at $p\in\{\bar{V}-2\bar{m}\bar{r}>0\}$ if and only if the matrix $h$ defined in (3.58) is positive definite at $p$. Alternatively, one can compare (3.59) with (3.50) to reach the same conclusion.
Let ${\mycal N}\approx\mathbb{T}^{N}$ denote any principal orbit of the action of the isometry group generated by the Killing vectors fields $\partial_{\phi^{i}}$, and consider the pullback, which we denote by $h_{\mycal N}$, of $h=h_{ij}d\phi^{i}d\phi^{j}$ to $\mycal N$}. When non-degenerate, $h_{\mycal N}$ is the metric induced by $\bar{g}$ on $\mycal N$}. Since $h_{\mycal N}$ is Riemannian on the set $\{\bar{V}-2\bar{m}\bar{r}<0\}$, which is stably causal as $\bar{\nabla}\bar{r}$ is timelike there, we have proved:
@@
1. All connected components of the set $\{\bar{V}-2\bar{m}\bar{r}<0\}$ are stably causal, with $\bar{\nabla}\bar{r}$ timelike there.
@@
and then $\bar{\nabla}\bar{t}$ is timelike there.
@@ 3.9.1 Negative Ricci curvature
$\ -\frac{\ell^{-2}(1-\ell^{-2}\bar{r}^{2})}{\bar{W}}\bigg[|\bar{\mu}|^{4}\big(\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}n^{i}dn^{i}}{\bar{\Xi}_{i}}\big)^{2}+2|\bar{\mu}|^{3}(\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(n^{i})^{2}d|\bar{\mu}|}{\bar{\Xi}_{i}})(\sum_{j=1}^{N}\frac{\bar{a}_{j}^{2}n^{j}dn^{j}}{\bar{\Xi}_{i}})\bigg]$
@@ 3.10 Black hole regions
We claim that $({\mycal M}_{\mathrm{KS}},\bar{g})$ contains a black hole region or a white hole region, depending upon the choice of time orientation. To see this, recall that the gradient of $\bar{r}$ is a time function on the set $\{\bar{V}-2\bar{m}\bar{r}<0\}\cap\{\bar{r}<\bar{r}_{H}\}$. Let us choose the time orientation so that $\bar{r}$ is a time function on $\{\bar{V}-2\bar{m}\bar{r}<0\}$ near $\bar{r}_{H}$, thus $\bar{r}$ is strictly decreasing along every causal curve in this last region. Hence no causal curve that intersects
@@ 4 Projection diagrams
$\breve{g}:=\$$\bar{W}(1-\ell^{-2}\bar{r}^{2})d\bar{t}^{2}+\frac{\bar{U}d\bar{r}^{2}}{\bar{V}-2\bar{m}\bar{r}}+\frac{2\bar{m}\bar{r}}{\bar{U}}\biggl(\bar{W}d\bar{t}+\sum_{i=1}^{N}\frac{\bar{a}_{i}(\bar{\mu}^{i})^{2}d\phi^{i}}{\bar{\Xi}_{i}}\biggr)^{2}$
@@
with $\pm N^{2}$ given by (3.50), thus $\breve{\gamma}$ is obtained by ignoring the $h$-part of $\hat{g}$. It follows from (4.2) and (3.47) that, in the region where $h$ is positive definite we have, for any vector $X$,
@@
In order to extract out of $\mathring{\gamma}\equiv\breve{U}^{-1}\breve{\gamma}$ a two-dimensional Lorentzian metric which depends only upon $\bar{r}$, we define
@@ 4.1 $\bar{V}-2\bar{m}\bar{r}<0$
$\pm N^{2}=\$$-\frac{\bar{W}(\bar{V}-2\bar{m}\bar{r})}{\left(1+\frac{2\bar{m}\bar{r}}{\bar{U}}\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{(\bar{r}^{2}+\bar{a}_{i}^{2})\bar{\Xi}_{i}}\right)\prod_{i=1}^{N}(\bar{r}^{2}+\bar{a}_{i}^{2})}\,$, (4.9)
@@
$\mathring{\gamma}(X,X)=\$$\breve{U}^{-1}N^{2}(X^{\bar{t}})^{2}+\frac{\bar{F}^{2}}{\bar{V}-2\bar{m}\bar{r}}(X^{\bar{r}})^{2}$
$\geq\$$|\bar{V}-2\bar{m}\bar{r}|\bar{H}_{-}(X^{\bar{t}})^{2}-\frac{\bar{F}^{2}}{|\bar{V}-2\bar{m}\bar{r}|}(X^{\bar{r}})^{2}$
@@ 4.2 $\bar{V}-2\bar{m}\bar{r}>0$
On this region the coordinate $\bar{r}$ has spacelike gradient, which implies that the tensor field $\gamma$ of (3.47) has Lorentzian signature. On the subset where the matrix $h$ of (3.58) is positive definite we have $\pm N^{2}<0$, hence we need to choose the negative sign in the left-hand side of (4.9). Then
$\mathring{\gamma}(X,X)=\$$-\breve{U}^{-1}N^{2}(X^{\bar{t}})^{2}+\frac{\bar{F}^{2}}{\bar{V}-2\bar{m}\bar{r}}(X^{\bar{r}})^{2}$
$\geq\$$-(\bar{V}-2\bar{m}\bar{r})\bar{H}_{+}(X^{\bar{t}})^{2}+\frac{\bar{F}^{2}}{\bar{V}-2\bar{m}\bar{r}}(X^{\bar{r}})^{2}$
@@ 4.3 Further extensions
Figure 4.1: The projection diagram for an extended spacetime with $\bar{m}>0$, in which there are no closed causal curves, with $\bar{V}-2\bar{m}\bar{r}$ having two zeros $\ell<\bar{r}_{-}<\ell<\bar{r}_{+}$. Both horizons are of bifurcate-type, with non-zero surface gravity. The light blue regions are where $\nabla\bar{r}$ is timelike and the pink ones where $\nabla\bar{t}$ is timelike. The solid-blue curves represent the level sets of constant $\bar{r}$ while the dashed-red those of constant $\bar{t}$. The Kerr–Schild extension is indicated by the regions marked I-III. The diagram can be continued indefinitely in the vertical direction so that a shift of the figure by two diamonds generates a $\mathbbm{Z}$-group of isometries. Distinct maximal analytic extensions can be obtained by quotienting the spacetime by distinct subgroups of $\mathbbm{Z}$, leading to closed timelike curves through every point. Further distinct maximal analytic extensions can be obtained by removing, e.g., one or more bifurcation surfaces and passing to covering spaces.
Figure 4.2: The projection diagram for an extended spacetime with $\bar{m}>0$ in which there are no closed causal curves and $\bar{V}-2\bar{m}\bar{r}$ has three zeros $\ell<\bar{r}_{1}<\bar{r}_{2}<\ell<\bar{r}_{3}$, with the zero at $\bar{r}_{2}$ of order two. The coloring of diamonds and the dotting of curves are as in Figure 4.1. The extension can be continued in all directions. The horizons at $\bar{r}_{2}$ are degenerate, in the sense that they have vanishing surface gravity $\kappa$, the other ones have non-zero $\kappa$.
Figure 4.3: The projection diagram for an extended spacetime with $\bar{m}>0$ in which there are no closed causal curves and $\bar{V}-2\bar{m}\bar{r}$ has four zeros $\ell<\bar{r}_{1}<\bar{r}_{2}<\bar{r}_{3}<\ell<\bar{r}_{4}$. The extension can be continued to the whole Minkowskian plane in all directions. All horizons are of bifurcate-type, with non-zero surface gravity. The nature of the diamonds and of the curves are as in Figure 4.1. The Kerr–Schild extension is the union of the regions marked I to V. The extension is symmetric with respect to reflections across vertical and horizontal lines passing through the bifurcation points of the Killing horizons. Shifts of the plane by two diamonds along the vertical axis and by two diamonds along the horizontal axis generate a $\mathbbm{Z}^{2}$ group of isometries of the depicted spacetime when extended to the whole of $\mathbb{R}^{2}$; taking quotients by distinct subgroups of $\mathbbm{Z}^{2}$ leads to distinct, non-isometric, analytic extensions of the original metric, possibly but not necessarily introducing causality violations. Further distinct extensions can be obtained by passing to distinct coverings of the projection plane.
@@
Figure 4.4: Projection diagrams for the Kerr–Newman anti-de Sitter metrics with two distinct zeros of the Boyer-Lindquist function $\Delta_{r}$ (left diagram) and one double zero (right diagram), from COS , generated using SzybkaOelz . The arrows represent the orientation of the orbits of the Killing vector $\partial_{t}$. The shaded area is the projection of the Carter time-machine region, including the singular ring, and should be removed from the diagram for a faithful representation of causality. An identical diagram is obtained for our metrics in the two-zeros case of Figure 4.1 when causality violations occur with $\bar{m}>0$.
@@ 5 Odd dimensions? Positive $\Lambda$?
$t$ $=i\bar{t}\,,\quad r=-i\bar{r}\,,\quad\,m=i^{-(d-1)}\bar{m}\,,\quad a_{i}=-i\bar{a}_{i}\,,\quad 1\leq i<N\,$$\qquad$ (5.1)
@@ Appendix A Determinant
$h\equiv h_{ij}d\phi^{i}d\phi^{j}=\$$\underbrace{\sum_{i=1}^{N}\frac{\bar{r}^{2}+\bar{a}_{i}^{2}}{\bar{\Xi}_{i}}(\bar{\mu}^{i})^{2}(d\phi^{i})^{2}}_{A}+\frac{2\bar{m}\bar{r}}{\bar{U}}\underbrace{\biggl(\sum_{i=1}^{N}\frac{\bar{a}_{i}(\bar{\mu}^{i})^{2}d\phi^{i}}{\bar{\Xi}_{i}}\biggr)^{2}}_{uu^{T}}\,$, (A.7)
$N^{j}=\$$h^{ij}\frac{2\bar{m}\bar{r}\bar{W}\bar{a}_{i}(\bar{\mu}^{i})^{2}}{\bar{U}\bar{\Xi}_{i}}\,$, (A.8)
$\pm N^{2}=\$$\bar{W}(1-\ell^{-2}\bar{r}^{2})+\frac{2\bar{m}\bar{r}}{\bar{U}}\bar{W}^{2}-h_{ij}N^{i}N^{j}\,.$ (A.9)
Next, there is the standard decomposition of the determinant of $\hat{g}$ in terms of $\pm N^{2}$ and $h$;
@@
Now to calculate $\pm N^{2}$ we need to know $N^{i}$ and hence we need also $h^{ij}$. Assuming from now on that we are away from zeros of $\left(1+\frac{2\bar{m}\bar{r}}{\bar{U}}\,u^{T}A^{-1}u\right)$, using (A.2) we obtain
@@
Now we can obtain $\pm N^{2}$ from (A.9);
$\pm N^{2}$ $=\bar{W}(1-\ell^{-2}\bar{r}^{2})+\frac{2\bar{m}\bar{r}}{\bar{U}}\bar{W}^{2}\left(1-\frac{2\bar{m}\bar{r}}{\bar{U}}\frac{u^{T}A^{-1}u}{1+\frac{2\bar{m}\bar{r}}{\bar{U}}\,u^{T}A^{-1}u}\right)$
@@
$\pm N^{2}$ $=\frac{\bar{W}}{1+\frac{2\bar{m}\bar{r}}{\bar{U}}\,u^{T}A^{-1}u}\left((1-\ell^{-2}\bar{r}^{2})+\frac{2\bar{m}\bar{r}}{\bar{U}}\left[1+\sum_{i=1}^{N}\frac{\bar{a}_{i}^{2}(\bar{\mu}^{i})^{2}}{\bar{r}^{2}+\bar{a}_{i}^{2}}\right]\right)$
@@
Note that we have recovered (A) without dividing by $1-\ell^{-2}\bar{r}^{2}$. However, in this alternative calculation we have instead a division-with-cancellation by $\left(1+\frac{2\bar{m}\bar{r}}{\bar{U}}\,u^{T}A^{-1}u\right)$, which vanishes at the boundary of the causality-violating regions. Since the previous calculation did not necessitate such a division, and since the zero-sets do not overlap in the fully-rotating case by (3.39), we conclude that (A.18) holds without the restrictions mentioned. Hence:
@@
Steps 2-4 are taken care of since they only set up the problem. In particular, since we did not change the coordinates $\bar{\mu}^{i}$ for $1\leq i\leq N$ we have simply $\tilde{\gamma}_{ij}=\bar{g}_{ij}$. Thus, the only thing left to calculate is
@@
To calculate the inverse metric we plug (A.2) into (A.28) to obtain (note that the expression $1+u\cdot D^{-1}\cdot u$ appearing below has already been calculated in (A))